HUBBLE SPACE TELESCOPE DAILY REPORT #2046 PERIOD COVERED: 0000Z (UTC) 01/23/98 - 0000Z (UTC) 01/26/98 Daily Status Report as of 026/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Six Sets of NIC/1/2/3 7235 (Observations of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Hubble Deep Field area HDF-123645. This program observes a portion of the Hubble Deep Field with all three NICMOS cameras in the F160W and F110W filters to produce infrared images equivalent in depth to the existing WF/PC-2 HDF images. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Eight Sets of STIS/MA1/MA2 7920 (STIS MAMA Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make parallel field, ultraviolet imaging to search for Lyman-break galaxies at z>2.0, and z>1.0. The observations were completed as planned, and no problems were reported. 1.4 Completed STIS/CCD 7900 (STIS CCD Pure Parallels for NIC/3 HDF) The Space Telescope Imaging Spectrograph (CCD) was used to make pure parallel, ultraviolet, slitless, observations with the NIC/3 Camera. This proposal will run during the orbits devoted to NICMOS HDF prime proposal 7235. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of clusters HBC-362, GL0388 and GL-406. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed Eight Sets of NIC/1/3 7813 (NIC/3 HDF Darks) The Near Infrared Camera (Cameras One and Three) was used to make internal dark calibration images which will be used when processing the Hubble Deep Field images during the first NICMOS Camera Three campaign. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxy IC4710 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.8 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at its nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed NIC/3 7815 (NICMOS Earth Flats) The Near Infrared Camera (Camera Three) was used to make a Earth flat field calibration image. This proposal obtains a series of earth flats for a subset of narrow and medium band NIC/3 filters in advance of the NIC/3 campaign. These flats will be used to generate the master flats for narrow-band filters and to determine the OTA illumination pattern as a function of wavelength. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Three Sets of NIC/1/2/3 7456 (Water in Protoplanetary Disks) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the protoplanetary disk around VSSG-17, RY-Tau and DN-Tau to search for water in their disks. This survey will be the first to take a detailed look at the gas phase properties of pre-main-sequence disks at AU distances. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 23 Successful : 23 Scheduled Reacquisitions : 23 Successful : 23 2.2 FHST Updates: Scheduled : 46 Successful : 46 2.3 Operations Notes: The Battery #2 Capacity Check was completed nominally at 024/1929Z. A STIS Status Buffer Message was received at 024/203032Z indicating that a STIS cal lamp retry was experienced. ROP NS-12 was performed at 024/2120Z. HSTAR 6327 documents this event. The NICMOS-3 Hubble Deep Field campaign continues into this week. For the entire weekend period, all HDF proposals executed nominally while the associated guide star performance had no anomalous activity. The NICMOS instrument is performing well. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH